OpenStudy (anonymous):

In Martin’s class of 20 students, 4 students report that social studies is their favorite class and 75% of the students report that PE is their favorite class. Show your work to explain how to determine whether more students prefer social studies or PE.

OpenStudy (anonymous):

@xXStudierxX

OpenStudy (anonymous):

Do you have any answer choices?

OpenStudy (anonymous):

nope

OpenStudy (anonymous):

75% is 3/4 of 20. 20-4=16, which means there are 16 students that prefer PE. 16 is greater than 4.

OpenStudy (studygurl14):

\(\tt\color{blue}{Percentage~of~people~who~say~social~studies~is~their~favorite~class:}\) \(\Large\frac{4}{20}\rightarrow\frac{1}{5}\rightarrow20\%\)

OpenStudy (studygurl14):

What is the higher percentage? \(\large20\%\) or \(\large75\%\)

OpenStudy (anonymous):

75

OpenStudy (anonymous):

Correct!

OpenStudy (studygurl14):

Your way works too @xXStudierxX . \(\Large\ddot\smile\)

OpenStudy (studygurl14):

Right. So that means more people say PE is their favorite subject than social studies.

OpenStudy (anonymous):

Thanks! StudyGurl14

OpenStudy (studygurl14):

:)